OpenStudy (anonymous):

What quadratic function is formed with a directrix of y=-2 and a focus of (1, 6)?

OpenStudy (cybershadow):

see the vertex must be exactly between focus and directrix i.e 1 and -2

OpenStudy (cybershadow):

eh nvm, Ur directrix is y = -2 , so it wud be between, 6 and -2

OpenStudy (anonymous):

How would I turn that into a function?

OpenStudy (cybershadow):

function wud be (x-h)^2 = 4(a)(x-k)^2

OpenStudy (cybershadow):

where h and k are coordinates of vertex

OpenStudy (cybershadow):

*(x-h)^2 = 4 (a) (y-k)^2 it shud be, nevermind the typo

OpenStudy (anonymous):

I'm really confused. :/

OpenStudy (cybershadow):

thats the equation of parabola , even i m confused over how to find the values of vertex :S ,once we get the value of vertex, we can find it

OpenStudy (cybershadow):

give me a minute

OpenStudy (anonymous):

okay. .-.

OpenStudy (cybershadow):

ok as the directrix is y = -2 and focus is (1,6) , the vertex must be midway between these two

OpenStudy (cybershadow):

that is (1,2)

OpenStudy (cybershadow):

got it so far?

OpenStudy (anonymous):

I think so.

OpenStudy (cybershadow):

ok now equation of parabola with vertex(h,k) i.e. (1,2) in our case is (y-h)^2 = 4a(x-k)^2 where a is distance between vertex and directrix

OpenStudy (cybershadow):

that makes our equation (y-2)^2 = 4a(x-1) ^2

OpenStudy (cybershadow):

all clear so far?

OpenStudy (anonymous):

It needs to be in the y= form though o:

OpenStudy (cybershadow):

whats a quadratic equation?

OpenStudy (cybershadow):

we can solve it and get it in y = x^2 + constant form

OpenStudy (anonymous):

Ohh

OpenStudy (cybershadow):

got it?

OpenStudy (anonymous):

I think so.

OpenStudy (cybershadow):

rite ,then lets find a , i.e. distance between vertex and directrix, and how wud we do that?

OpenStudy (anonymous):

So the vertex is 1,2 and the directrix is 0, -2?

OpenStudy (cybershadow):

yerr

OpenStudy (anonymous):

is the distance sqrt(17)?

OpenStudy (cybershadow):

rite , approx 4 then

OpenStudy (cybershadow):

(y-2)^2 = 4a(x-1) lets use this equation now

OpenStudy (anonymous):

Would the answer to my question by any chance be f(x)=1/8(x-1)^2-2?

OpenStudy (cybershadow):

ehh i guess i messed up with the parabola equation

OpenStudy (anonymous):

Now I'm confused.. again xD

OpenStudy (cybershadow):

no, its correct lol

OpenStudy (anonymous):

Is that the right answer? o.o

OpenStudy (cybershadow):

wait lemme solve

OpenStudy (anonymous):

or would it be f(x)=-1/8(x+1)^2-2? o.o

OpenStudy (cybershadow):

the equation shud be (x-1)^2 = 4(y-2)

OpenStudy (cybershadow):

thats where i messed up

OpenStudy (cybershadow):

x^2+1-2x = 4x4(y-8)

OpenStudy (cybershadow):

ah why did i oopen it, We just need to find the y lol

OpenStudy (anonymous):

wait wait. would this be a parabola that's got a smile or a frown? xD (i don't remember the exact word for it.. oops.)

OpenStudy (cybershadow):

see the vertex 's coordinates

OpenStudy (cybershadow):

they are (1,2)

OpenStudy (anonymous):

f(x)=1/16(x-1)^2+2 has a vertex of 1,2 o:

OpenStudy (cybershadow):

now can ya tell? :P

OpenStudy (anonymous):

would it be f(x)=1/16(x-1)^2+2? o.o

OpenStudy (cybershadow):

yes, !!

OpenStudy (anonymous):

:DD yaya! thank you!

OpenStudy (cybershadow):

if they got same vertex, means both are same parabolas

OpenStudy (cybershadow):

rite, DOne! lol sorry took me time to recall all the geometry things, havent studied it in a while

OpenStudy (anonymous):

it's fine. xD thanks though! :)

OpenStudy (cybershadow):

you're welcome :)